OpenStudy (anonymous):

the x- intercept of a line is -5 and the y-intercept of the line is -2. what is the equation of the line?

OpenStudy (anonymous):

Help please :)

OpenStudy (anonymous):

@iLoveMath0101

OpenStudy (anonymous):

ok lets see.

OpenStudy (anonymous):

Ax + By = C do yu know that formula

OpenStudy (anonymous):

No, that's why I'm having such a hard time. I'm great in Science though :)

OpenStudy (anonymous):

Oh ok, A(x) + B(y)=C A= x intercept B= y intercept

OpenStudy (anonymous):

Is your answers multiple choice?

OpenStudy (anonymous):

Here are the possible solutions..A.Y=-5/2 x-5 B.Y=2/5x+2 C.Y=5/2 x-2 D.Y=-2/5x-2

OpenStudy (anonymous):

I got it

OpenStudy (anonymous):

One second

OpenStudy (anonymous):

Its rise over run It would become \[y=\frac{ 2 }{ -5 }-2\]

OpenStudy (anonymous):

You are awesome!!!!!!!!

OpenStudy (anonymous):

Welcome!

OpenStudy (anonymous):

so D then ?

OpenStudy (anonymous):

yep